3x^2-2x+1=0
solve the equation and present in solution set notation

Answer by nerdybill(6962)

3x^2-2x+1=0
.
Factoring the left side:
(3x+1)(x-1) = 0
.
Setting each term on the left to zero to find the set:
3x+1 = 0
3x = -1
x = -1/3
.
x-1 = 0
x = 1
.
Solution set:
x = {-1/3, 1}